suppose that good X is perfectly divisible (you can buy non-integer quantities).

The supply curve is given by Qs=1.2P - 1.

The demand curve is given by Qd=38 - 1.1P.

the equilibrium quantity is?

The equilibrium price is?

answer these 2 questions.
Answer #1

At the equilibrium price the demand and supply are equal.

we can write it as Qs= QD

or 1.2P -1 = 38 -1.1P

1.2P + 1.1P = 38+1

2.3P = 39

P = 16.95 the equilibrium price in the market is 16.95.

Quantity will be 38 - 1.1 P

38 - 1.1 (18.64)

38 - 20.50.

= 17.5 the equilibrium quantity will be 17.5